OpenStudy (anonymous):

S.O.S the perimeter of a sign is 84 inches, and the length is 6 inches longer than width. Width and length are both needed.

OpenStudy (anonymous):

p=2l+2w 84=2(6+w)+2w 84=12+2w+2w 76=4w w=19 l=6+w l=6+19=25

OpenStudy (anonymous):

I Don't think this is right? I get 18, 24

OpenStudy (anonymous):

and you found out where my math goes awry occasionally. sorry! But for any other perimeter problem you may have, start it out the same way or a similar way, cool?